Vector A has a magnitude of 6.0 m and points 30 south of east. Vector B has a magnitude of 4.0 m and points 30° west of south. The resultant vector A B is given by
Answer #1

Magnitude of the vector A = 6m

Magnitude of the vector B = 4m

Angle between the vectors A and B is (30+30)=60°

Now the sum of the vectors A + B =?

(A+B)--A2 B2 2 * A * B * cos(60)

(A + B)2-62 + 42 + 2 * 6 * 4 * cos(60)

(A B) 3616 +48 1/2

(A + B)2-52 + 24 = 76

(A B) (76)/2 8.717m